# The Complexity of HyperQPTL

### v2

@@ 3.2 HyperQPTL Satisfiability
  • For third-order variables $\mathcal{Y}_{j}$ and second-order variables $Y$,
    ${\mathit{hyp}}(Y\in\mathcal{Y}_{j})=(\texttt{m}_{j})_{\pi_{Y}})$,
  • For second-order variables $Y$ and first-order variables $y$,
    ${\mathit{hyp}}(y\in Y)=\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\texttt{x}_{\pi_{Y}})$,
  • For first-order variables $y,y^{\prime}$, ${\mathit{hyp}}(y<y^{\prime})=\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{F}\vphantom{a}}\nolimits\texttt{x}_{\pi_{y^{\prime}}})$,
  • For first-order variables $y_{1},y_{2},y$, ${\mathit{hyp}}(y_{1}+y_{2}=y)=\exists\pi.\ \texttt{add}_{\pi}\land\newline \mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{1}}}\land\texttt{arg1}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{2}}}\land\texttt{arg2}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\texttt{res}_{\pi})$.
  • For first-order variables $y_{1},y_{2},y$, ${\mathit{hyp}}(y_{1}\cdot y_{2}=y)=\exists\pi.\ \texttt{mult}_{\pi}\land\newline \mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{1}}}\land\texttt{arg1}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{2}}}\land\texttt{arg2}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\texttt{res}_{\pi})$.
@@
An induction shows that we indeed have that $(\mathbb{N},+,\cdot,<,\in)\models\varphi$ if and only if $f(\varphi)$ is satisfiable: Over the structure of $\psi$, one proves for all variable assignments $\alpha$ of the free variables of a subformula $\psi^{\prime}$ that $(\mathbb{N},+,\cdot,<,\in),\alpha\models\psi^{\prime}$ if and only if $T_{\alpha},\Pi_{\alpha},0\models{\mathit{hyp}}(\psi^{\prime})$, where $T_{\alpha}$ is a set containing enough traces to satisfy $\theta_{\mathit{all}}\wedge\theta_{(+,\cdot)}$, which are additionally marked according to the values $\alpha(\mathcal{Y}_{j})$ such that $\theta_{\mathit{cons}}$ is satisfied, and where $\Pi_{\alpha}$ mimics the assignment $\alpha$ to first- and second-order variables as explained above. Then, the remaining existential third-order quantifiers are mimicked by the choice of $T$, as also explained above. We leave the fully formal definition and the actual inductive proof, which follows immediately from the remarks above, to the reader.
Now, consider a language $L$ in $\Sigma^{2}_{1}$, i.e., it is of the form $\{n\in\mathbb{N}\mid(\mathbb{N},+,\cdot,<,\in)\models\varphi(n)\}$ where $\varphi(x)$ is a formula of the form $\exists\mathcal{Y}_{1}.\ \ldots\exists\mathcal{Y}_{k}.\ \psi(x,\mathcal{Y}_{1},\ldots,\mathcal{Y}_{k})$, with first-order variable $x$ and third-order variables $\mathcal{Y}_{j}$, and with $\psi$ only containing first-order and second-order quantifiers.
There is a function that, given $n\in\mathbb{N}$ returns a first-order formula $\theta_{=n}(x)$ with a single free first-order variable $x$ that is only satisfied in (,+,,<,)($\mathbb{N}$$\cdot,$$\in$) if $x$ is assigned $n$. Furthermore, this function can be implemented in polynomial time in $\log n$. Hence, we have $n\in L$ if and only if $f(\exists x.\ \theta_{=n}(x)\wedge\varphi(x))$ is satisfiable. Thus, we have completed the desired reduction: $\mathrm{Hyper\-QPTL}$ satisfiability is indeed $\Sigma^{2}_{1}$-hard. ∎
Now, we consider the upper bound. Here, we use the fact that traces can be encoded as sets of natural numbers via an encoding that is “implementable” in arithmetic. This will allow us to encode the satisfiability problem in arithmetic.
@@
Formulas $\varphi$ of $\mathrm{Hyper\-QPTL}$ can be encoded via natural numbers $\mathit{enc}(\varphi)$ in a way that the encoding is implementable in first-order arithmetic. Relying on such an encoding, we present a formula $\theta(x)$ of arithmetic with a single free (first-order) variable $x$ such that a $\mathrm{Hyper\-QPTL}$ sentence $\varphi$ is satisfiable if and only if $(\mathbb{N},+,\cdot,<,\in)\models\theta(\mathit{enc}(\varphi))$. To obtain our desired result, $\theta$ may only use existential third-order quantification, but can use arbitrary second- and first-order quantification. Intuitively, we will express the existence of a nonempty set of traces and the existence of some structures that prove that this set is a model of $\varphi$, generalizing the technique used to prove $\Sigma_{1}^{1}$-membership of $\mathrm{Hyper\-LTL}$ satisfiability [7]. Both the set of traces and the additional structures can be encoded as third-order objects. For the remainder of the proof, let us fix a $\mathrm{Hyper\-QPTL}$ sentence $\varphi$ over $\mathrm{AP}$ and let $\psi$ be its maximal quantifier-free subformula. To simplify our constructions, we assume w.l.o.g. that each trace variable and each proposition is quantified at most once in $\varphi$ and that $\mathrm{AP}$ and the set of trace variables appearing in $\varphi$ are disjoint finite subsets of $\mathbb{N}$. All these properties can always be achieved by renaming the trace variables and the propositions.
@@
In the semantics of $\mathrm{Hyper\-QPTL}$ as defined above, propositional quantification updates the set of traces the formula is evaluated over (the $T$ in $T,\Pi,i\models\varphi$) by uniformly updating the truth values of a proposition in the traces. Here, however, it is more convenient to leave the $T$ unchanged and instead keep track of the truth values assigned to propositions using an assignment mapping propositions q to sequences in $(2^{\{\texttt{q}\}})^{\omega}$. Hence, in this proof we will be working with two assignments, a trace assignment and a propositional assignment. To clearly distinguish them, we use the symbols $\Pi_{t}$ and $\Pi_{p}$, respectively.
Note that $\Pi_{t}$ and $\Pi_{p}$ contain enough information to evaluate quantifier-free formulas: For a position $i$ we define
  • $\Pi_{t},\Pi_{p},i\models\texttt{p}_{\pi}$ if $\texttt{p}\in\Pi_{t}(\pi)(i)$,
  • $\Pi_{t},\Pi_{p},i\models\texttt{q}$ if $\texttt{q}\in\Pi_{p}(\texttt{q})(i)$,
  • $\Pi_{t},\Pi_{p},i\models\lnot\psi^{\prime}$ if $\Pi_{t},\Pi_{p},i\not\models\psi^{\prime}$,
  • $\Pi_{t},\Pi_{p},i\models\psi_{1}^{\prime}\lor\psi_{2}^{\prime}$ if $\Pi_{t},\Pi_{p},i\models\psi_{1}^{\prime}$ or $\Pi_{t},\Pi_{p},i\models\psi_{2}^{\prime}$,
  • $\Pi_{t},\Pi_{p},i\models\mathop{\mathbf{X}\vphantom{a}}\nolimits\psi^{\prime}$ if $\Pi_{t},\Pi_{p},i+1\models\psi^{\prime}$, and
@@
  • $\Pi_{t},\Pi_{p},i\models\mathop{\mathbf{F}\vphantom{a}}\nolimits\psi^{\prime}$ if there is a $i^{\prime}\geq i$ such that $\Pi_{t},\Pi_{p},i^{\prime}\models\psi^{\prime}$.
@@
To capture quantification of traces and propositions, we use Skolem functions for the existentially quantified variables. Fix some nonempty set $T$ of traces over $\mathrm{AP}$. Let $\pi$ be an existentially quantified trace variable in $\varphi$ and let $\pi_{1},\ldots,\pi_{k}$ be the trace variables that are universally quantified before $\pi$ and let $\texttt{q}_{1},\ldots,\texttt{q}_{\ell}$ be the propositional variables that are universally quantified before $\pi$. Then, a $T$-Skolem function for $\pi$ is a function mapping a $k$-tuple of traces in $T$ and an $\ell$-tuple of traces (the $i$-th one over $\{\texttt{q}_{i}\}$) to a trace in $T$. Similarly, let q be an existentially quantified propositional variable in $\varphi$ and let $\pi_{1},\ldots,\pi_{k}$ be the trace variables that are universally quantified before $\pi$ and let $\texttt{q}_{1},\ldots,\texttt{q}_{\ell}$ be the propositional variables that are universally quantified before $\pi$. Then, a $T$-Skolem function for q is a function mapping a $k$-tuple of traces in $T$ and an $\ell$-tuple of traces (the $i$-th one over $\{\texttt{q}_{i}\}$) to a trace over $\{\texttt{q}\}$.
@@
Consider a pair $(\Pi_{t},\Pi_{p})$ of a trace assignment and a propositional assignment such that all universally quantified trace variables in $\varphi$ are mapped to some trace in $T$ by $\Pi_{t}$. We say that $(\Pi_{t},\Pi_{p})$ is consistent with a $T$-Skolem function $f$ for an existentially quantified trace variable $\pi$ if $\Pi(\pi)=f(\Pi_{t}(\pi_{1}),\ldots,\Pi_{t}(\pi_{k}),\Pi_{p}(\texttt{q}_{1}),\ldots,\Pi_{p}(\texttt{q}_{\ell}))$ and it is consistent with a $T$-Skolem function $f$ for an existentially quantified propositional variable q if $\Pi(\texttt{q})=f(\Pi_{t}(\pi_{1}),\ldots,\Pi_{t}(\pi_{k}),\Pi_{p}(\texttt{q}_{1}),\ldots,\Pi_{p}(\texttt{q}_{\ell}))$. In both cases, the $\pi_{j}$ and the $\texttt{q}_{j}$ are the trace and propositional variables universally quantified before $\pi$ and q, respectively.
@@
Now, we have $T\models\varphi$ (in the semantics introduced in Subsection 3.1) if and only if there are $T$-Skolem functions for all existentially quantified trace variables and $T$-Skolem functions for all existentially quantified propositional variables such that
@@
  • for every propositional assignment $\Pi_{p}$ whose domain contains all propositional variables in $\psi$, and that is consistent with the Skolem functions,
@@
In fact, $e_{\psi,\Pi_{t},\Pi_{p}}$ is completely characterized by the following consistency conditions, which only depend on $\Pi_{t}$ and $\Pi_{p}$:
  • $e_{\psi,\Pi_{t},\Pi_{p}}(\texttt{p}_{\pi},i)=1$ if and only if $\texttt{p}\in\Pi_{t}(\pi)(i)$,
@@
Note that all these conditions are expressible in first-order arithmetic.
Now, the formula $\theta$ expresses the existence of the following objects:
@@
  • A function $\mathcal{E}\colon(2^{\mathbb{N}})^{*}\times\mathbb{N}\times\mathbb{N}\rightarrow\{0,1\}$ to be interpreted as expansions as follows: if the first argument $A$ encodes a trace assignment $\Pi_{t}$ and a propositional assignment $\Pi_{p}$, then $x,y\mapsto\mathcal{E}(A,x,y)$ encodes the expansion $e_{\psi,\Pi_{t},\Pi_{p}}$, i.e., $x$ encodes a subformula of $\psi$ and $y$ a position.
Note that $\mathcal{S}$ and $\mathcal{E}$ can also be encoded as sets of sets of natural numbers, using the fact that there is a bijection between $(2^{\mathbb{N}})^{*}$ and $2^{\mathbb{N}}$ using the pairing function, and that a function can be encoded by its graph, i.e., a function from a set $A$ to a set $B$ can be encoded as as a subset of $A\times B$. All these encodings and bijections are implementable in arithmetic.
@@
  • $\mathcal{T}$ does indeed only contain encodings of traces over $\mathrm{AP}$, i.e., it encodes a set $T$ of traces.
  • The image of $\mathcal{S}$ only contains traces over $T$, if the first argument is a trace variable, and traces over $\{\texttt{q}\}$, if the first input is a propositional variable q.
  • For every $A\in(2^{\mathbb{N}})^{*}$ encoding a trace assignment $\Pi_{t}$ and a propositional assignment $\Pi_{p}$ such that all universally quantified trace variables in $\varphi$ are mapped by $\Pi_{t}$ to some trace encoding of a trace in $T$ and all universally quantified propositional variables q in $\varphi$ are mapped by $\Pi_{p}$ to an encoding of a trace over $\{\texttt{q}\}$, $\mathcal{E}(A,\cdot,\cdot)$ satisfies the consistency conditions (i.e., $\mathcal{E}$ indeed encodes the expansion) and we have $\mathcal{E}(A,n_{0},0)=1$, where $n_{0}$ is the natural number encoding $\psi$.
We leave the tedious, but standard, details to the reader. ∎

### v3

@@ 3.2 HyperQPTL Satisfiability
  • For third-order variables $\mathcal{Y}_{j}$ and second-order variables $Y$, ${\mathit{hyp}}(Y\in\mathcal{Y}_{j})=(\texttt{m}_{j})_{\pi_{Y}}$,
  • For second-order variables $Y$ and first-order variables $y$, ${\mathit{hyp}}(y\in Y)=\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\texttt{x}_{\pi_{Y}})$,
  • For first-order variables $y,y^{\prime}$, ${\mathit{hyp}}(y<y^{\prime})=$
    $\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{F}\vphantom{a}}\nolimits\texttt{x}_{\pi_{y^{\prime}}})$,
  • For first-order variables $y_{1},y_{2},y$, ${\mathit{hyp}}(y_{1}+y_{2}=y)=\exists\pi.\ \texttt{add}_{\pi}\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{1}}}\land\texttt{arg1}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{2}}}\land\texttt{arg2}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\texttt{res}_{\pi})$ and ${\mathit{hyp}}(y_{1}\cdot y_{2}=y)=\exists\pi.\ \texttt{mult}_{\pi}\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{1}}}\land\texttt{arg1}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y_{2}}}\land\texttt{arg2}_{\pi})\land\mathop{\mathbf{F}\vphantom{a}}\nolimits(\texttt{x}_{\pi_{y}}\land\texttt{res}_{\pi})$.
@@
An induction shows that we have that $(\mathbb{N},+,\cdot,<,\in)\models\varphi$ if and only if $f(\varphi)$ is satisfiable: Over the structure of $\psi$, one proves for all variable assignments $\alpha$ of the free variables of a subformula $\psi^{\prime}$ that $(\mathbb{N},+,\cdot,<,\in),\alpha\models\psi^{\prime}$ if and only if $T_{\alpha},\Pi_{\alpha},0\models{\mathit{hyp}}(\psi^{\prime})$, where $T_{\alpha}$ is a set containing enough traces to satisfy $\theta_{\mathit{\!all}}\wedge\theta_{(+,\cdot)}$, which are additionally marked according to the values $\alpha(\mathcal{Y}_{j})$ such that $\theta_{\mathit{cons}}$ is satisfied, and where $\Pi_{\alpha}$ mimics the assignment $\alpha$ to first- and second-order variables as explained above. Then, the remaining existential third-order quantifiers are mimicked by the choice of $T$, as also explained above. We leave the fully formal definition and the actual inductive proof, which follows from the remarks above, to the reader.
Now, consider a language $L$ in $\Sigma^{2}_{1}$, i.e., it is of the form $\{n\in\mathbb{N}\mid(\mathbb{N},+,\cdot,<,\in)\models\varphi(n)\}$ where $\varphi(x)$ is a formula of the form $\exists\mathcal{Y}_{1}\ldots\exists\mathcal{Y}_{k}.\ \psi(x,\mathcal{Y}_{1},\ldots,\mathcal{Y}_{k})$, with first-order variable $x$ and third-order variables $\mathcal{Y}_{j}$, and with $\psi$ only containing first-order and second-order quantifiers.
There is a function that, given $n\in\mathbb{N}$ returns a first-order formula $\theta_{=n}(x)$ with a single free first-order variable $x$ that is only satisfied in (,+,,<,)($\mathbb{N}$$\cdot,$$\in$) if $x$ is assigned $n$. This function can be implemented in polynomial time in $\log n$. So, we have $n\in L$ if and only if $f(\exists x.\ \theta_{=n}(x)\wedge\varphi(x))$ is satisfiable. Thus, we have completed the desired reduction: $\mathrm{Hyper\-QPTL}$ satisfiability is indeed $\Sigma^{2}_{1}$-hard. ∎
Now, we consider the upper bound. Here, we use the fact that traces can be encoded as sets of natural numbers via an encoding that is “implementable” in arithmetic. This allows us to encode the satisfiability problem in arithmetic.
@@
Formulas $\varphi$ of $\mathrm{Hyper\-QPTL}$ can be encoded via natural numbers $\mathit{enc}(\varphi)$ in a way that the encoding is implementable in first-order arithmetic (e.g., by a Gödel numbering). Relying on such an encoding, we present a formula $\theta(x)$ of arithmetic with a single free (first-order) variable $x$ such that a $\mathrm{Hyper\-QPTL}$ sentence $\varphi$ is satisfiable if and only if $(\mathbb{N},+,\cdot,<,\in)\models\theta(\mathit{enc}(\varphi))$. To obtain our result, $\theta$ may only use existential third-order quantification, but arbitrary second- and first-order quantification.
Intuitively, we express the existence of a nonempty set of traces and the existence of some structures that prove that this set is a model of $\varphi$, generalizing the technique used to prove $\Sigma_{1}^{1}$-membership of $\mathrm{Hyper\-LTL}$ satisfiability [7]. These structures include in particular Skolem functions for the existential variables. Then, we express in arithmetic that every trace assignment (which can be encoded by a type $1$ object) that is consistent with the Skolem functions satisfies the maximal quantifier-free subformula of $\varphi$. To evaluate this subformula, we implement the semantics of quantifier-free $\mathrm{Hyper\-QPTL}$ in arithmetic by existentially quantifying a function that assigns to each trace assignment $\Pi$ and each quantifier-free subformula $\psi^{\prime}$ of $\varphi$ the truth value of $\psi^{\prime}$ with respect to $\Pi$. This function is again an existentially quantified type $2$ object and its correctness can be expressed in first-order arithmetic. In the following, we introduce the necessary encoding of traces, an equivalent semantics of $\mathrm{Hyper\-QPTL}$ that is more convenient for the construction here, Skolem functions, and then the construction of the formula $\theta$.
For the remainder of the proof, let us fix a $\mathrm{Hyper\-QPTL}$ sentence $\varphi$ over $\mathrm{AP}$. To simplify our constructions, we assume w.l.o.g. that each trace variable and each proposition is quantified at most once in $\varphi$. Also, we assume w.l.o.g. that $\mathrm{AP}$ and the set of trace variables appearing in $\varphi$ are disjoint finite subsets of $\mathbb{N}$. These properties can be guaranteed by renaming trace variables and propositions.
@@
Recall that our plan is to existentially quantify a model (a set of traces), which is encoded by a set of sets of natural numbers, i.e., a type $2$ object. Then, we want to evaluate the formula over that model using arithmetic. Furthermore, the semantics of $\mathrm{Hyper\-QPTL}$ as defined in Subsection 3.1 updates the model $T$ every time a proposition is quantified, i.e., from $T$ to $T[\texttt{q}\mapsto t_{\texttt{q}}]$ for some trace $t_{\texttt{q}}\in(2^{\texttt{q}})^{\omega}$. To capture this update naively, in case q is universally quantified, requires universal quantification of a type $2$ object that captures the encoding of $T[\texttt{q}\mapsto t_{\texttt{q}}]$. However, this would not yield the desired $\Sigma_{1}^{2}$ upper bound.
@@
However, we do not need to update the model $T$ as long as we keep track of $t_{\texttt{q}}$, as all traces in $T$ have the same truth values w.r.t. q, i.e., those of $t_{\texttt{q}}$. Hence, we keep track of the truth values assigned to quantified propositions using an assignment mapping propositions q to sequences in $(2^{\{\texttt{q}\}})^{\omega}$. Thus, in this proof we work with two assignments, a trace assignment and a propositional assignment. To clearly distinguish them, we use the symbols $\Pi_{t}$ and $\Pi_{p}$, respectively.
Next, we introduce the modified semantics. For a position $i$ and $T$, $\Pi_{t}$, and $\Pi_{p}$ as above, we define
  • $T,\Pi_{t},\Pi_{p},i\models\texttt{p}_{\pi}$ if p is in the domain of $\Pi_{p}$ and $\texttt{p}\in\Pi_{p}(\texttt{p})(i)$, or if p is not in the domain of $\Pi_{p}$ and $\texttt{p}\in\Pi_{t}(\pi)(i)$,
  • $T,\Pi_{t},\Pi_{p},i\models\texttt{q}$ if $\texttt{q}\in\Pi_{p}(\texttt{q})(i)$,
  • $T,\Pi_{t},\Pi_{p},i\models\lnot\psi^{\prime}$ if $T,\Pi_{t},\Pi_{p},i\not\models\psi^{\prime}$,
  • $T,\Pi_{t},\Pi_{p},i\models\psi_{1}^{\prime}\lor\psi_{2}^{\prime}$ if $T,\Pi_{t},\Pi_{p},i\models\psi_{1}^{\prime}$ or $T,\Pi_{t},\Pi_{p},i\models\psi_{2}^{\prime}$,
  • $T,\Pi_{t},\Pi_{p},i\models\mathop{\mathbf{X}\vphantom{a}}\nolimits\psi^{\prime}$ if $T,\Pi_{t},\Pi_{p},i+1\models\psi^{\prime}$,
@@
  • $T,\Pi_{t},\Pi_{p},i\models\mathop{\mathbf{F}\vphantom{a}}\nolimits\psi^{\prime}$ if there is an $i^{\prime}\geq i$ such that $T,\Pi_{t},\Pi_{p},i^{\prime}\models\psi^{\prime}$,
  • $T,\Pi_{t},\Pi_{p},i\models\exists\pi.\,\varphi$ if there exists a trace $t\in T$ such that $T,\Pi_{t}[\pi\mapsto t],\Pi_{p},i\models\varphi$,
  • $T,\Pi_{t},\Pi_{p},i\models\forall\pi.\ \varphi$ if for all traces $t\in T$ we have $T,\Pi_{t}[\pi\mapsto t],\Pi_{p},i\models\varphi$,
  • $T,\Pi_{t},\Pi_{p},i\models\exists\texttt{q}.\ \varphi$ if there exists a trace $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$ such that $T,\Pi_{t},\Pi_{p}[\texttt{q}\mapsto t_{\texttt{q}}],i\models\varphi$, and
  • $T,\Pi_{t},\Pi_{p},i\models\forall\texttt{q}.\ \varphi$ if for all traces $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$ we have $T,\Pi_{t},\Pi_{p}[\texttt{q}\mapsto t_{\texttt{q}}],i\models\varphi$.
We have that $T,\Pi_{\emptyset},0\models\varphi$ (i.e., in the semantics as in Subsection 3.1) if and only if $T,\Pi_{t,\emptyset},\Pi_{p,\emptyset},0\models\varphi$ (i.e., in the new semantics) for all nonempty $T$ and all $\mathrm{Hyper\-QPTL}$ sentences $\varphi$, where $\Pi_{t,\emptyset}$ and $\Pi_{p,\emptyset}$ denote the assignments with empty domain. Also, for quantifier-free formulas $\psi^{\prime}$, $T,\Pi_{t},\Pi_{p},i\models\psi^{\prime}$ is independent of $T$, i.e., we have $T,\Pi_{t},\Pi_{p},i\models\psi^{\prime}$ if and only if $T^{\prime},\Pi_{t},\Pi_{p},i\models\psi^{\prime}$ for all $T$ and $T^{\prime}$. This is due to the fact that $T$ is only referred to in the cases of trace quantification in the new semantics. This is useful later when we construct $\theta$.
To simplify the construction of the formula $\theta(x)$, we capture existential quantification of traces and propositions by Skolem functions. This allows us replace all quantifiers by a single universal quantifier that ranges over assignments that are consistent with the Skolem functions.
@@
Fix some nonempty set $T$ of traces over $\mathrm{AP}$. Let $\pi$ be an existentially quantified trace variable in $\varphi$ and let $\pi_{1},\ldots,\pi_{k}$ be the trace variables that are universally quantified before $\pi$ and let $\texttt{q}_{1},\ldots,\texttt{q}_{\ell}$ be the propositional variables that are universally quantified before $\pi$. A $T$-Skolem function for $\pi$ is a function mapping a $k$-tuple of traces in $T$ and an $\ell$-tuple of traces (the $i$-th one over $\{\texttt{q}_{i}\}$) to a trace in $T$. Similarly, let q be an existentially quantified propositional variable in $\varphi$ and let $\pi_{1},\ldots,\pi_{k}$ be the trace variables that are universally quantified before $\pi$ and let $\texttt{q}_{1},\ldots,\texttt{q}_{\ell}$ be the propositional variables that are universally quantified before $\pi$. A $T$-Skolem function for q is a function mapping a $k$-tuple of traces in $T$ and an $\ell$-tuple of traces (the $i$-th one over $\{\texttt{q}_{i}\}$) to a trace over $\{\texttt{q}\}$.
@@
Consider a pair $(\Pi_{t},\Pi_{p})$ of a trace assignment and a propositional assignment such that all universally quantified trace variables in $\varphi$ are mapped to some trace in $T$ by $\Pi_{t}$ and all universally quantified propositional variables q are mapped to some trace in $(2^{\{\texttt{q}\}})^{\omega}$ by $\Pi_{p}$. We say that $(\Pi_{t},\Pi_{p})$ is consistent with a $T$-Skolem function $f$ for an existentially quantified trace variable $\pi$ if $\Pi_{t}(\pi)=f(\Pi_{t}(\pi_{1}),\ldots,\Pi_{t}(\pi_{k}),\Pi_{p}(\texttt{q}_{1}),\ldots,\Pi_{p}(\texttt{q}_{\ell}))$ and it is consistent with a $T$-Skolem function $f$ for an existentially quantified propositional variable q if $\Pi_{t}(\texttt{q})=f(\Pi_{t}(\pi_{1}),\ldots,\Pi_{t}(\pi_{k}),\Pi_{p}(\texttt{q}_{1}),\ldots,\Pi_{p}(\texttt{q}_{\ell}))$. In both cases, the $\pi_{j}$ and the $\texttt{q}_{j}$ are the trace and propositional variables universally quantified before $\pi$ and q, respectively.
@@
Recall that we have fixed a $\mathrm{Hyper\-QPTL}$ sentence $\varphi$. Let $\psi$ be the maximal quantifier-free subformula of $\varphi$. Now, we have $T\models\varphi$ (in the semantics from Subsection 3.1) if and only if there are $T$-Skolem functions for all existentially quantified trace variables and for all existentially quantified propositional variables such that
@@
  • for every propositional assignment $\Pi_{p}$ whose domain contains all propositional variables in $\psi$, that maps universally quantified q to traces over $\{\texttt{q}\}$, and that is consistent with the Skolem functions,
@@
In fact, $e_{\psi,\Pi_{t},\Pi_{p}}$ is completely characterized by the following consistency conditions that only depend on $\Pi_{t}$ and $\Pi_{p}$:
  • $e_{\psi,\Pi_{t},\Pi_{p}}(\texttt{p}_{\pi},i)=1$ if and only if either p is in the domain of $\Pi_{p}$ and $\texttt{p}\in\Pi_{p}(\texttt{p})(i)$ or if p is not in the domain of $\Pi_{p}$ and $\texttt{p}\in\Pi_{t}(\pi)(i)$,
@@
Note that all these conditions are expressible using a formula of first-order arithmetic (that depends on $\psi$).
Now, the formula $\theta$ expresses the existence of the following third-order objects:
@@
  • A function $\mathcal{E}\colon(2^{\mathbb{N}})^{*}\times\mathbb{N}\times\mathbb{N}\rightarrow\{0,1\}$ to be interpreted as expansions as follows: if the first argument $A$ encodes a trace assignment $\Pi_{t}$ and a propositional assignment $\Pi_{p}$ whose domains contain exactly the variables quantified in $\varphi$, then $x,y\mapsto\mathcal{E}(A,x,y)$ encodes the expansion $e_{\psi,\Pi_{t},\Pi_{p}}$, i.e., $x$ encodes a subformula of $\psi$ and $y$ a position.
Note that $\mathcal{S}$ and $\mathcal{E}$ can also be encoded as sets of sets of natural numbers, using the fact that a sequence $(S_{1},\cdots,S_{k})\in(2^{\mathbb{N}})^{*}$ can be encoded by the set $\{\mathit{pair}(n,j)\mid n\in S_{j}\}\in 2^{\mathbb{N}}$, and that a function $f\colon\mathbb{N}\rightarrow\mathbb{N}$ can be encoded by $\{\mathit{pair}(n,f(n))\mid n\in\mathbb{N}\}\subseteq\mathbb{N}$. All these encodings are implementable in arithmetic.
@@
  • $\mathcal{T}$ does indeed only contain encodings of traces over $\mathrm{AP}$, i.e., it encodes a set $T$ of traces (using the formula $\varphi_{\mathit{isTrace}}$ from above).
  • The image of $\mathcal{S}$ only contains traces over $T$, if the first argument is a trace variable, and traces over $\{\texttt{q}\}$, if the first input is a propositional variable q (using a formula $\varphi_{\mathit{format}}(\mathcal{S},\mathit{enc}(\varphi))$ of second-order arithmetic).
  • For every $A\in(2^{\mathbb{N}})^{*}$ encoding a trace assignment $\Pi_{t}$ and a propositional assignment $\Pi_{p}$ such that
    • $(\Pi_{t},\Pi_{p})$ is consistent with the Skolem functions encoded by $\mathcal{S}$,
    • all universally quantified trace variables in $\varphi$ are mapped by $\Pi_{t}$ to some trace encoding of a trace in $T$, and
    • all universally quantified propositional variables q in $\varphi$ are mapped by $\Pi_{p}$ to an encoding of a trace over $\{\texttt{q}\}$,
    we require that $\mathcal{E}(A,\cdot,\cdot)$ satisfies the consistency conditions (i.e., $\mathcal{E}$ indeed encodes the expansion) and that we have $\mathcal{E}(A,n_{0},0)=1$, where $n_{0}$ is the natural number encoding $\psi$ (using a formula $\varphi_{\mathit{correct}}(\mathcal{S},\mathit{enc}(\varphi))$ of second-order arithmetic).
Thus, $\theta$ has the form
$\exists\mathcal{T}.\ \exists\mathcal{S}.\ \exists\mathcal{E}.\ \exists Y.\ Y\in\mathcal{T}\wedge\forall Y.\ (Y\in\mathcal{T}\rightarrow\varphi_{\mathit{isTrace}}(Y))\wedge\varphi_{\mathit{format}}(\mathcal{S},\mathit{enc}(\varphi))\wedge\varphi_{\mathit{correct}}(\mathcal{S},\mathit{enc}(\varphi)).$
We leave the tedious, but standard, construction of $\varphi_{\mathit{format}}(\mathcal{S},\mathit{enc}(\varphi))$ and $\varphi_{\mathit{correct}}(\mathcal{S},\mathit{enc}(\varphi))$ to the reader.
Putting all the pieces together yields that $\varphi$ is satisfiable if and only if $(\mathbb{N},+,\cdot,<,\in)\models\theta(\mathit{enc}(\varphi))$. ∎